A doorbell needs 0.5A and 8.0V and is therefore connected to a transformer secondary coil. The primary coil has 3000 turns and is connected to a 230V wall socket (a) how many turns does the secondary coil need? (b) How big is the current in the primary coil?
1
Expert's answer
2019-05-15T09:37:32-0400

Transformation ratio of idealized transformer:


"U_1\/U_2=w_1\/w2"



Turns secondary coil:

"w_2=w_1*U_2\/U_1=3000*8\/230=104 turns""U_1*I_1=\/(U_2*I_2)=P"

Current in the primary coil:

"I_1=U_2*I_2\/U_1=8*0.5\/230=0.017A"

Answer: 104 turns, 17mA